6. LO3 A ball is thrown vertically upward with a speed of 25m/s. What is maximum height reached by the ball? a. 1.25m b. 31m c. 50 m d. 0.5 m
Answer #1

using 3rd equation of motion

v^2 = 2 g h

25^2 = 2*9.8* h

h = 31.89 m

========

do comment in case any doubt, will reply for sure.. Goodluck